WHAT PERCENTAGE OF 4 IS 5?
Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

WHAT PERCENTAGE OF 4 IS 5?
----------
(5/4)*100%

Answer by fractalier(6550)

Change this into math...
x/100 * 4 = 5
x/100 = 5/4
x = 500/4 = 125%